# Observational Signatures of a Previous Dynamical Instability in Multi-planet M-Dwarf Systems

### v1

@@ Abstract
We identify observational signatures suggesting a history of dynamical instability in 26 out of 34 M-dwarf multi-planet systems containing no large planets. These systems may have primarily formed in a gas-rich environment, potentially hosted more planets and were more compact. We extend previous simulations of the formation of the TRAPPIST-1 system to $100\,\rm Myr$ to test the stability of these systems without gas. We find the absence of a strong mean motion resonance in the innermost planet pair and the absence of three body resonances throughout the system are likely to result in the merging and ejection of planets after the gas disk disperses. The runs that experience such an instability tend to produce final systems with lower multiplicities, period ratios larger than two, increased orbital spacings, higher planetary angular momentum deficits, and slightly smaller mass ratios between adjacent planets. Remarkably, we find these same trends in the observations of M-dwarf multi-planet systems containing no large planets. Our work allows us to identify specific systems that may have experienced an instability and suggests that only $\sim 25\%$ of these systems formed in their current observed state while most systems were likely more compact and multiplicitous earlier in time. Previous research indicates that systems that have experienced a late stage giant impact may host planets potentially more habitable than the systems that did not. With this in mind, we suggest systems around M-dwarfs that contain period ratios larger than two be given priority in the search for habitable worlds.
@@ 5.1 Methodology
We note that 53% of the planets in the Big PR class were observed by radial velocity (RV) and the remaining 47% were observed by transit whereas in the small PR class, 76% of planets were observed by transit, 14% by RV and the remaining 10% were observed either by transit timing variations or orbital brightness modulation. We note these observation methods as each method has its own associated observational biases. Most notably, the majority of the small PR planets in our sample were observed by transit which is biased against long period orbits. Thus, it is possible that these systems also harbor longer period planets that could change the trends we describe here.
Additionally, observational biases may make the observed multiplicities incomplete. This incompleteness is likely not significant for the Small PR group as most of these planets have been detected via transit and the presence of additional planets could be detected by transit timing variations. On the other hand, since most of the planets in the Big PR group were observed via RV, it is possible these observations are missing additional planets.
@@ 6 Discussion and Conclusions
Admittedly, we are unable to differentiate between systems that formed primarily through pebble accretion and went unstable and systems that formed primarily through late stage core accretion. Still, in both scenarios the planets would likely have experienced late stage giant impacts which should have implications for their final properties, unlike the systems that form primarily through pebble accretion but do not experience any instabilities after the gas disk dissipates. For example, giant impacts can produce a moon system (Canup & Asphaug, 2001; Canup, 2004), internal heating that can result in plate tectonics and volcanism (Yuan et al., 2024), the generation of a secondary atmosphere (Liggins et al., 2022), a molten core (Brian Tonks & Jay Melosh, 1992) and in turn, potentially, a magnetic field (Bullard, 1949). These are all reminiscent of Earth and are thought to be beneficial to the habitability of a planet. Thus, we suggest that systems around M-dwarfs that contain larger period ratios and lower planet multiplicities be given priority in the search for habitable worlds.

### v2

@@ Abstract
We identify observational signatures suggesting a history of dynamical instability in 26 out of 34 M-dwarf multi-planet systems containing no large planets. These systems may have primarily formed in a gas-rich environment, potentially hosted more planets and were more compact. We extend previous simulations of the formation of the TRAPPIST-1 system to $100\,\rm Myr$ to test the stability of these systems without gas. We find the absence of a strong mean motion resonance in the innermost planet pair and the absence of three body resonances throughout the system are likely to result in the merging and ejection of planets after the gas disk disperses. The runs that experience such an instability tend to produce final systems with lower multiplicities, period ratios larger than two, increased orbital spacings, higher planetary angular momentum deficits, and slightly smaller mass ratios between adjacent planets. Remarkably, we find these same trends in the observations of M-dwarf multi-planet systems containing no large planets. Our work allows us to identify specific systems that may have experienced an instability and suggests that only $\sim 25\%$ of these systems formed in their current observed state while most systems were likely more compact and multiplicitous earlier in time. Previous research indicates that planets that have experienced a late stage giant impact may potentially be more habitable than those that did not. With this in mind, we suggest systems around M-dwarfs that contain period ratios larger than two be given priority in the search for habitable worlds.
@@ 5.1 Methodology
We note that 53% of the planets in the Big PR class were observed by radial velocity (RV) and the remaining 47% were observed by transit whereas in the small PR class, 76% of planets were observed by transit, 14% by RV and the remaining 10% were observed either by transit timing variations or orbital brightness modulation. We note these observation methods as each method has its own associated observational biases. Most notably, the majority of the small PR planets in our sample were observed by transit which is biased against long period orbits however, the presence of additional planets could be detected by transit timing variations. On the other hand, since most of the planets in the Big PR group were observed via RV, it is possible these observations are missing additional small planets that could change the trends we describe here.
@@ 6 Discussion and Conclusions
Admittedly, we are unable to differentiate between systems that formed primarily through pebble accretion and went unstable and systems that formed primarily through late stage core accretion. Still, in both scenarios the planets would likely have experienced late stage giant impacts which should have implications for their final properties, unlike the systems that form primarily through pebble accretion but do not experience any instabilities after the gas disk dissipates. For example, giant impacts can produce a moon system (Canup & Asphaug, 2001; Canup, 2004), internal heating that can result in plate tectonics and volcanism (Yuan et al., 2024), the generation of a secondary atmosphere (Liggins et al., 2022), a molten core (Brian Tonks & Jay Melosh, 1992) and in turn, potentially, a magnetic field (Bullard, 1949). These are all reminiscent of Earth and are thought to be beneficial to the habitability of a planet. Thus, we suggest that systems around M-dwarfs that contain period ratios greater than two be given priority in the search for habitable worlds.